OpenStudy (anonymous):

Identify the number of solutions and its type for the equation 9x2 − 12x + 4 = 0.

alones (alones):

Hey Welcome to Openstudy :) you have any multiple choice dear?

alones (alones):

9x^2−12x+4=0 Step 1: Factor left side of equation. (3x−2)(3x−2)=0 Step 2: Set factors equal to 0. 3x−2=0 or 3x−2=0 x=2/3 @ParthKohli am I right?